factor completely:
X^4-2X^2-16X +32 = 0

Answer by rapaljer(4671)

I LOVE factoring, but I think you copied this one wrong!!! If I am wrong, please send me an Email (or a Thank You note!), and I'll try again for you.

This is a "Factoring by Grouping" problem where you must group the first two and the second two together!

x^3 - 2x^2 - 16x + 32 = 0

x^2(x-2) - 16(x-2) = 0

Now, take out the common factor of (x-2).
(x-2) (x^2 - 16) = 0

Now, factor the difference of squares:
(X-2)(x-4)(x+4) = 0

Finally, this is an EQUATION, so solve the equation by setting each factor equal to zero:
x= 2, x=4, x=-4

I have a LOT of explanations, examples, and FULL COLOR solutions to problems like this on my own website. I even have VIDEOS of me teaching this type of factoring in my younger days before I retired and my hair was still dark!! This is a GREAT chapter if you get a good explanation of it--not nearly as hard as your textbook is making it.

My website is all FREE with NO ADVERTISING, and I'm not selling anything! I just want to make this as easy as I can for you.

To see my website go to www.mathinlivingcolor.com, named after my theme "Math in Living Color." Look for a single link near the bottom of this page that will take you to my Homepage. Once on my Homepage, look for "Basic, Intermediate and College Algebra: One Step at a Time", and select "Basic Algebra". Look in "Chapter 2" for the appropriate topics!

As for the videos, look for the link on my Homepage called "Rapalje Videos in Living Color", and choose either Basic or Intermediate Algebra. Choose the videos that have to do with Factoring.

If you have questions on other topics, just send me an Email! I'll be glad to help you find the right page on my website. I hope you like my website and that you will recommend it to family and friends! Remember, my goal is to make math EASIER than your own textbook is making it, and it's ALL FREE!!!


Dr. Robert J. Rapalje, Retired
Seminole State College of Florida

www.mathinlivingcolor.com.